OpenStudy (mathmath333):

How to identify if a probability question applies the concept of binomial distribution ?

OpenStudy (michele_laino):

for example, if we have to study this event: "I toss 25 coins, and I want to know what is the probability to get, for example, 10 heads"

OpenStudy (michele_laino):

in such example, I make 25 trials, and I want to get 10 successes. the probability to get a head is 1/2, the probability to ge a tail is also 1/2, and the requested probability, is given by this quantity: \[\Large P = \left( {\begin{array}{*{20}{c}} {25} \\ {10} \end{array}} \right){\left( {\frac{1}{2}} \right)^{10}}{\left( {\frac{1}{2}} \right)^{\left( {25 - 10} \right)}}\]

OpenStudy (mathmath333):

Is the following question using binomial distribution ? The bag \(A\) contains \(3\) black balls and \(5\) red balls while bag \(B\) contains \(4\) black balls and \(7\) red balls , If the \(2\) balls extracted simultaneosuly from any of the two bags are found to be black, find the probability that they are found from bag \(B\).

Parth (parthkohli):

Probably conditional probability. Don't ask me though, I haven't studied it.

Miracrown (miracrown):

Well, for your question, it is a binomial case when we are looking for the number of success in a sample. I mean, it could be (similar to the exercise) the number of red balls selected from a bag.

OpenStudy (mathmath333):

ok thnks

Miracrown (miracrown):

yw :)

OpenStudy (mathmath333):

I will also request @Michele_Laino to comment .

OpenStudy (sparrow2):